PLEAE SHOW WORK

A.What is the expected return of the following portfolio?13.5%

Stock

E(R)

Investment

A

10%

$50,000

B

15%

30,000

C

20%

20,000

B. nWhat is the expected return, beta, and standard deviation of the following portfolio? βp = 0.975   

Security

Weight

βi

E(Ri)

σi

ρA,B

A

25%

1.5

10%

40%

0.4

B

75%

0.8

20%

30%

C. The following annual returns for Stock E are projected over the next year for three possible states of the economy. What is the stock’s standard deviation, variance, and mean of returns?E(R) = 8.5% ; σ = 22.70% ; mean = $7.50 ; standard deviation = $2.50

State

Prob

E(R)

Boom

10%

40%

Normal

60%

20%

Recession

30%

- 25%
Answer #1

a) E(R) = wa x ra + wb x rb + wc x rc = (50,000 x 10% + 30,000 x 15% + 20,000 x 20%) / (50,000 + 30,000 + 20,000) = 13.50%

b) Expected Returns = 25% x 10% + 75% x 20% = 17.5%

Beta of the portfolio = 25% x 1.5 + 75% x 0.8 = 0.975

Std. Dev. = [(25% x 40%)^2 + (75% x 30%)^2 + (2 x 25% x 75% x 40% x 30% x 0.4]^(1/2) = 28.04%

c) E(R) = 10% x 40% + 60% x 20% + 30% x -25% = 8.5%